OpenStudy (anonymous):

cos(90+A)=-sinA. verify that it is an identity. I know that cos(90-A)=sinA but I am drawing a blank on how to get from there to a verification.

OpenStudy (accessdenied):

You know that: cos (90 - A) = sin A What happens when we trade A for -A in that equation?

OpenStudy (anonymous):

cos(90+A)=-sinA is that all though?

OpenStudy (anonymous):

ohh then you could do -sinA=-sinA as a substitution right?

OpenStudy (accessdenied):

I believe that is all we need. Since cos(90 - A) = sin A should be true for any angle A. We can choose B=-A such that cos(90 - B) = sin B cos (90 - (-A)) = sin (-A). Simplify, and use the fact that sine is odd. sin(-A) = - sinA cos(90 + A) = -sin A And all we had done was change A for -A, so it still works for all values of A.

OpenStudy (anonymous):

thank you so much!